OpenStudy (anonymous):

Integrate the following: INT e^2x cos3xdx

OpenStudy (anonymous):

I'll teach you a new technique. Its called complexifying the integral. First you must know that: \[e^{ix}=\cos(x)+i \sin(x)\] Where if I said: \[\Re[e^{ix}]=\cos(x) ; \Im[e^{ix}]=\sin(x)\] With that being said observe that: \[e^{2x+3ix}=e^{2x}e^{3\xi}=e^{2x}(\cos(3x)+i \sin(3x)); \Re=e^{2x}\cos(3x)\] So that means the integral must be the real part of the complex number above^. So the integral is the same as: \[\Re \left[ \int\limits e^{(2+3i)x} dx \right]=\Re \left[ \frac{e^{(2+3i)x}}{2+3i} \right]\] Multiply the top and bottom by the complex conjugate and expand out the e^(whatever) This gives: \[\Re \left[ \frac{(e^{2x}\cos(3x)+e^{2x}\sin(3x))(2-3i)}{(2+3i)(2-3i)} \right]\]

OpenStudy (anonymous):

From here multiply out and find the real part. Multiplying gives: \[\Re \left[ \frac{1}{13} [ e^{2x}(2\cos(3x)+\sin(3x))+e^{2x}(2\sin(3x)-3\cos(3x))] \right]=\frac{e^{2x}}{13}(2\cos(3x)+\sin(3x))+C\]

OpenStudy (anonymous):

After sin(3x) on the last line there should be a +C.

OpenStudy (anonymous):

I hope this isn't too "complex" -.-

OpenStudy (anonymous):

Also, there should be a 3 in front of the sin(3x) in the solution. I just mistyped again.

OpenStudy (anonymous):

What do you think outkast?